Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the cost of goods completed and transferred out during April
using the FIFO process costing method.
Answer #1

Cost of goods completed and transferred out = $498,536

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion cost Material Conversion cost
Beginning WIP             18,000 38% 71% 62% 29%
Units introduced             80,000
Total units to be accounted for             98,000
Completed and Transferred unit             54,000 0% 0% 100% 100%
Ending WIP             44,000 0% 0% 28% 74%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion cost
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP             18,000 62%                   11,160 29%                          5,220
From units started/Introduced             36,000 100%                   36,000 100%                        36,000
Total             54,000                   47,160                        41,220
Total Units Material Conversion cost
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total             44,000 28%                   12,320 74%                        32,560
Total EUP             98,000                   59,480                        73,780

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion cost TOTAL
Cost incurred in Current Period $ 190,336 $ 232,407 $ 422,743
Total Equivalent Units                                              59,480                                                       73,780
Cost per Equivalent Units $ 3.20 $ 3.15 $ 6.35
Cost of Beginning WIP $ 217,781
Total cost to be accounted for $ 640,524

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred               54,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                  11,160 $ 3.20 $ 35,712
Conversion cost                     5,220 $ 3.15 $ 16,443 $ 52,155
B From units started & completed
Material                  36,000 $ 3.20 $ 115,200
Conversion cost                  36,000 $ 3.15 $ 113,400 $ 228,600
C=A+B Total $ 280,755
D Cost of Beginning WIP $ 217,781
E=C+D Total Cost of Units transferred $ 498,536
Cost of Ending WIP               44,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                  12,320 $ 3.20 $ 39,424
G Conversion cost                  32,560 $ 3.15 $ 102,564
H = F + G Total Cost of Ending WIP $ 141988.0
I = E + H Total Cost accounted for $ 640,524